At&T 10 year, 6% annual coupon rate, 1,000 par value bond is selling at $1,200. Interest on this bond is paid annually. Calculate both approximate yield maturity and exact yield to maturity.
Answer #1
Approximate Yield to maturity is calculated as follows:
Approximate Yield to maturity = Average income / Average investment
= $          40.00 / $    1,100.00
= 3.64%
Working:
Average income = (C+((F-P)/n)) Where,
= (60+((1000-1200)/10)) C = Annual coupon payment = 1000*6% = $                 60.00
= $          40.00 F = Maturity value = $           1,000.00
P = Current price = $           1,200.00
n = Time = 10
Average Investment = (F+P)/2
= (1000+1200)/2
= $    1,100.00
Exact Yield to maturity = =rate(nper,pmt,-pv,fv) Where,
= 3.58% nper = Time = 10
pmt = Coupon Payment = $              60.00
pv = Current price = $        1,200.00
fv = Maturity value = $        1,000.00